# LUSIFER: Language Universal Space Integration for Enhanced Multilingual Embeddings with Large Language Models

### v1

@@ 4 Experiment
In this section, we first introduce the benchmark datasets and evaluation metrics in Section 4.1. Then, we describe the experimental setup, including the model implementation, training data, and training details in Section 4.2. Afterward, we present the main results in Section 4.3, and analyze the effectiveness of LUSIFER’s components in Section 4.6. Finally, we visualize the LUSIFER’s representations in multilingual space to obtain insights into its lingual-agnostic capabilities in Section 4.7.
@@
Jina-embeddings-v3* Sturua et al. (2024) 59.84 61.23 62.88 58.94 66.74 78.35 58.51 64.71 73.57 64.96 64.19 61.54 68.96 49.20 63.83
mGTE-base* Zhang et al. (2024) 60.40 59.65 61.02 56.20 65.81 73.46 56.55 61.97 68.96 61.22 60.81 58.24 63.58 52.57 61.46
BGE-M3* Chen et al. (2024) 60.09 60.60 62.37 57.34 70.69 78.97 58.78 64.12 75.60 64.72 64.61 65.31 69.85 54.20 64.80
Multilingual-E5-large* Wang et al. (2024d) 61.91 61.97 62.91 59.40 71.30 78.08 55.21 63.41 76.53 66.55 63.75 63.67 67.32 51.55 64.54
UDEVER-Bloom-7B* Zhang et al. (2023) 55.83 56.39 59.73 54.38 64.32 68.70 48.97 55.02 67.60 58.54 55.96 55.13 61.00 47.41 57.78
SimCSE Gao et al. (2021b) 51.92 51.81 24.90 46.95 31.18 37.12 39.27 29.46 41.64 26.23 25.17 21.54 26.71 38.36 35.16
Contriever Izacard et al. (2022) 49.29 44.26 26.55 44.05 33.03 39.66 38.33 32.36 45.76 26.47 23.27 22.61 22.64 39.26 34.82
GTE-large Li et al. (2023) 62.29 51.66 33.49 50.13 38.88 44.67 43.07 30.27 51.98 27.02 20.38 22.97 22.75 41.40 38.64
BGE-en-1.5 Xiao et al. (2023) 63.27 51.65 32.79 50.84 38.50 49.73 43.28 30.81 51.16 31.11 25.28 26.34 23.02 41.96 39.98
E5-large Wang et al. (2024a) 60.12 52.41 26.81 51.00 37.99 39.47 43.86 31.32 53.59 28.84 24.57 23.48 22.03 43.25 38.48
ST5-XXL Ni et al. (2021c) 58.81 60.35 44.42 58.50 41.81 24.66 53.43 25.30 52.46 15.43 18.07 17.10 21.63 38.81 37.91
GTR-XXL Ni et al. (2021b) 58.12 54.39 41.94 53.21 37.96 24.67 50.08 25.14 53.88 15.23 17.35 15.92 22.12 40.57 36.47
E5-Mistral Wang et al. (2024b) 66.64 61.84 61.30 59.65 58.58 72.55 58.25 54.43 66.97 62.82 56.23 55.10 47.15 50.61 59.44
@@
Table 1: Comparative analysis of model performance across multiple languages and tasks. The table presents average metrics for each model, with the highest score for each language emphasized in bold. * denotes the models trained on extensive multilingual data.
@@
SimCSE Gao et al. (2021b) 7.41 18.35 39.71 37.95 0.18 20.72
Contriever Izacard et al. (2022) 9.75 22.94 34.55 41.72 0.03 21.80
GTE-large Li et al. (2023) 16.99 31.82 37.57 53.79 1.59 28.35
BGE-en-1.5 Xiao et al. (2023) 16.64 31.19 40.40 50.77 1.11 28.02
E5-large Wang et al. (2024a) 17.04 31.12 37.90 54.31 1.83 28.44
ST5-XXL Ni et al. (2021c) 20.82 41.68 56.19 59.02 1.76 35.89
GTR-XXL Ni et al. (2021b) 20.19 38.02 50.83 60.11 2.74 34.38
E5-Mistral Wang et al. (2024b) 31.54 54.75 81.12 71.37 21.92 52.14
@@
Table 2: Cross-lingual evaluation results. The table presents average metrics for each model over all languages of the datasets, with the highest score for each language emphasized in bold.
4.1 Benchmark
Figure 2 illustrates the tasks and datasets in our benchmark. Following Muennighoff et al. (2023), our benchmark includes five fundamental embedding tasks, with the evaluation protocol for each task adapted from the respective original papers. The benchmark involves 123 diverse datasets, including 48 Classification datasets, 24 Clustering datasets, 24 Retrieval datasets, 22 Semantic Textual Similarity STS datasets, and 5 Reranking datasets. The main metrics for each task are as follows: Classification: Accuracy, Clustering: V-measure Rosenberg and Hirschberg (2007), Retrieval: nDCG@10, STS: Pearson correlation based on cosine similarity Reimers et al. (2016), and Reranking: MAP. Following Lai et al. (2023), our benchmark covers 14 languages including 5 high-resource languages: English (en), Spanish (es), Russian (ru), French (fr), Vietnamese (vi); 6 medium-resource languages: Persian (fa), Indonesian (id), Arabic (ar), Finnish (fi), Korean (ko), Hindi (hi); 3 low-resource languages: Bengali (bn), Telugu (te), Swahili (sw).
Additionally, we evaluate models on cross-lingual retrieval tasks where the models need to perform text embedding tasks with queries and documents in different languages. These tasks feature 5 datasets, including Belebele Bandarkar et al. (2024), MLQA Lewis et al. (2020a), STS17, STS22 Agirre et al. (2016), and IndicCrosslingualSTS Ramesh et al. (2022), covering over 100 languages, including critically low-resource languages.
@@ 4.1 Benchmark
Figure 3: Performance comparison of LUSIFER and baseline models on Classification and Clustering tasks.
4.2 Experimental Setup
@@
Implementation Details. LUSIFER encompasses three key components: a multilingual encoder, a connector, and a target LLM. We employ XLM-R-large Conneau et al. (2020) as the multilingual encoder, Mistral-7B Jiang et al. (2023) as the English-centric target LLM, and a 2-layer feed-forward network with one trainable token as the connector. To facilitate efficient training, we leverage the LoRA framework Hu et al. (2022) for training of LUSIFER’s components. Furthermore, we employ GradCache Gao et al. (2021a), gradient checkpointing, mixed precision training, and FSDP Zhao et al. (2023) to minimize GPU memory requirements. The LUSIFER architecture and its training code are built on top of the Hugging Face Transformers Wolf et al. (2020) and Pytorch Lightning libraries Falcon and team (2024). We detail the training hyper-parameters for each stage in Table 4 of Appendix A.
Training Data. We only train LUSIFER on a diverse public English datasets. For alignment training, we use the combination of the English Wikipedia and questions-answering datasets. Specifically, we use subset of Wikitext-103 Merity et al. (2017) and MSMARCO Bajaj et al. (2018) for the masked reconstruction and autoregressive completion tasks, respectively. For representation finetuning, we adopt the retrieval datasets as follows: MS MARCO Bajaj et al. (2018), NQ Kwiatkowski et al. (2019), PAQ Lewis et al. (2021), HotpotQA Yang et al. (2018), SNLI Bowman et al. (2015), SQuAD Rajpurkar et al. (2016), ArguAna Wachsmuth et al. (2018), FiQA Maia et al. (2018) and FEVER Thorne et al. (2018). To address the lack of hard negatives in these datasets, we leverage an encoder-based model Wang et al. (2024a) to select the hard negatives on those datasets. Refer to Table 5 for the number of samples used in each dataset.
Baselines. We evaluate LUSIFER’s performance across the five fundamental embedding tasks on the benchmark datasets. We make comparisons with a variety of baseline models for embedding tasks which only trained/finetuned on mainly English data. Baselines include the following categories: dense retrieval models with Small Language Model (SLM) backbone: SimCSE Gao et al. (2021b), Contriever Izacard et al. (2022), GTE-large Li et al. (2023), BGE-en-1.5 Xiao et al. (2023), E5-large Wang et al. (2024a); and dense retrieval models with Large Language Model (LLM) backbone: GTR-XXL Ni et al. (2021b), ST5-XXL Ni et al. (2021c), E5-Mistral Wang et al. (2024b). Moreover, we include the following state-of-the-art multilingual embedding models which are trained on extensive multilingual data for reference: Jina-embeddings-v3 Sturua et al. (2024), mGTE-base Zhang et al. (2024), BGE-M3 Chen et al. (2024), Multilingual-E5-large Wang et al. (2024d), and UDEVER-Bloom-7B Zhang et al. (2023).
4.3 Main Results
Table 1 presents the main results of LUSIFER and baseline models on the benchmark datasets. LUSIFER achieves state-of-the-art performance in 10 out of 14 languages, with an average score of 62.63 across all languages, a 3.19 points improvement over the previous best-performing baseline, E5-Mistral (59.44) Wang et al. (2024b). Note that E5-Mistral is essentially the Mistral model fine-tuned on extensive proprietary synthetic data and supplemented with some multilingual data for training. Our results demonstrate that LUSIFER significantly enhances the multilingual capabilities of English-centric embedding LLM by aligning it with a multilingual encoder, enabling effective multilingual representation without requiring explicit multilingual training data. The improvements are particularly pronounced for medium and low-resource languages, with Telugu (te) showing the largest gain of 22.15 points over E5-Mistral. This highlights LUSIFER’s effectiveness in improving representation capabilities for traditionally underrepresented languages. Additionally, LUSIFER significantly outperforms the embedding models with SLM backbones, such as E5-large (38.48) and BGE-en-1.5 (39.98) which are trained on English data only, thus further demonstrating the benefits of combining multilingual encoder and LLM’s English-centric for text-embedding tasks in multilingual settings. Furthermore, even without explicit multilingual supervision, LUSIFER achieves competitive performance (62.63) compared to state-of-the-art multilingual models that require extensive multilingual training data, such as BGE-M3 (64.80) Chen et al. (2024) and Multilingual-E5-large (64.54) Wang et al. (2024d). These results further demonstrate the benefits of LUCIFER for multilingual representation learning while avoiding expensive multilingual data for text embeddings.
Baselines En Es Ru Fr Vi Fa Id Ar Fi Ko Hi Bn Te Sw Avg.
Table 3: Ablation study results of LUSIFER’s components. The table presents average metrics for each model, with the highest score for each language emphasized in bold.
4.4 Cross-Lingual Evaluation
Figure 4: Performance comparison of LUSIFER and baseline models on Reranking tasks.
Table 2 presents the results of LUSIFER and baseline models on the cross-lingual tasks. LUSIFER achieves the highest average score of 57.89, outperforming the previous best-performing baseline, E5-Mistral (52.14), by 5.75 points. Notably, LUSIFER demonstrates significant improvements in low-resource languages, as evidenced by its performance on the IndicCrosslingual dataset, where it achieves a score of 43.40, substantially higher than the next best baseline, E5-Mistral (21.92). These results underscore LUSIFER’s effectiveness in enhancing cross-lingual capabilities through efficient multilingual representation alignment, enabling the model to process text-embedding tasks across multiple languages effectively.
(a) Retrieval tasks
(b) STS tasks
Figure 5: Performance comparison of LUSIFER and baseline models on Retrieval and STS tasks.
4.5 Task-Specific Performance
Figure 3, 4, 5 present the performance comparison of LUSIFER and baseline models on Classification, Clustering, Reranking, Retrieval, and STS tasks. LUSIFER consistently outperforms the baseline models across 4 out of 5 tasks, with the largest improvements observed in Clustering and Retrieval tasks, especially in the medium and low-resource languages. However, the performance of LUSIFER in the Reranking tasks is slightly worse than the baseline models. This discrepancy may be attributed to the task’s complexity and the information loss in the alignment process between the multilingual encoder and the target LLM. Nevertheless, LUSIFER’s strong performance across a variety of tasks and languages highlights its ability to enhance multilingual representations without relying on explicit multilingual training data.
4.6 Ablation Study
To evaluate the effectiveness of LUSIFER’s components and training procerdure, we conduct an ablation study to analyze the impact of each component on the model’s performance. We compare the performance of LUSIFER with the following ablated versions: (1) LUSIFER with only finetuning connector in both alignment training and representation finetuning stages, (2) LUSIFER with frezzing the multilingual encoder while training the connector and the target LLM in both stages, (3) LUSIFER with only alignment training, i.e., alignment training without representation finetuning, (4) LUSIFER with only representation finetuning without alignment training. Table 3 presents the results of the ablation study. The full LUSIFER model achieves the highest average score of 62.63 across all languages, outperforming the ablated versions. Notably, the alignment training and representation finetuning stages both contribute to the model’s performance, with the representation finetuning stage showing a more substantial impact on the model’s performance. These results underscore the importance of each component in LUSIFER’s architecture and training process, highlighting the model’s effectiveness in enhancing multilingual representation capabilities.
4.7 Model Representation Visualization
(a) E5-Mistral (language)
(b) LUSIFER (language)
Figure 6: t-SNE representation of 200 randomly samples from the SIB200 dataset. The points are colored by the languages.
Figure 6 shows 2D scatter plots of representations from different models for 200 randomly sampled examples from the SIB200 dataset, visualized using t-SNE. The points are colored by the language of the samples. The t-SNE representation of E5-Mistral demonstrates a clearer separation between languages, with distinct clusters for each language. In contrast, the visualization of LUSIFER presents a more mixed distribution of languages, with overlapping clusters across different languages. This observation provides insights into LUSIFER’s lingual-agnostic capabilities, highlighting the model’s ability to bridge the gaps between representation spaces of different languages. These results suggest that LUSIFER’s alignment strategy enables the model to comprehend semantics across multiple languages effectively, facilitating zero-shot multilingual transfer. Overall, our experiments confirm the advantages of the representation alignment strategies in LUCIFER to effectively enable zero-shot multilingual transfer for LLM-based embedding methods.
5 Conclusion
In this work, we propose LUSIFER, a novel framework that enables effective multilingual representation without explicit multilingual training data. LUSIFER aligns a multilingual encoder with a target English-centric LLM through a minimal set of trainable parameters, facilitating zero-shot multilingual transfer. Our experimental results demonstrate that LUSIFER achieves state-of-the-art performance across diverse languages and tasks, outperforming existing baseline models. Moreover, LUSIFER significantly enhances cross-lingual capabilities, enabling the model to process text-embedding tasks across multiple languages effectively. Our work provides a promising direction for enhancing multilingual representation capabilities in English-centric embedding models, enabling global applicability without requiring extensive multilingual training data. In future work, we plan to explore additional alignment strategies and further investigate the impact of LUSIFER’s components on multilingual representation quality.
Appendix A Training Details
The code and dataset for training are available at https://github.com/hieum98/lusifer.
@@ Appendix A Training Details
Table 4: Training hyperparameters for each stage.
@@
Alignment Training Wikitext-103 Merity et al. (2017) 100,000
MSMARCO Bajaj et al. (2018) 100,000
Representation Finetuning MS MARCO Bajaj et al. (2018) 100,000
FEVER Thorne et al. (2018) 100,000
PAQ Lewis et al. (2021) 100,000
SNLI Bowman et al. (2015) 100,000
HotpotQA Yang et al. (2018) 97,800
SQuAD Rajpurkar et al. (2016) 97,400
FiQA Maia et al. (2018) 6,420
NQ Kwiatkowski et al. (2019) 3,420
ArguAna Wachsmuth et al. (2018) 1,280

### v2

@@ 4. Experiment
This section presents our experimental methodology and results. We first introduce benchmark datasets and evaluation metrics in Section 4.1, followed by our experimental setup including model implementation, training data, and procedures in Section 4.2. We then present our main findings in Section 4.3 and analyze LUSIFER’s cross-lingual capabilities in Section 4.4. We examine detail performance of LUSIFER across five embedding tasks in Section 4.5 and evaluate LUSIFER’s component effectiveness in Section 4.6. Finally, Section 4.7 visualizes LUSIFER’s representations in multilingual space to demonstrate its language-agnostic capabilities.
@@
Jina-embeddings-v3* (Sturua et al., 2024) 59.84 61.23 62.88 58.94 66.74 78.35 58.51 64.71 73.57 64.96 64.19 61.54 68.96 49.20 63.83
mGTE-base* (Zhang et al., 2024) 60.40 59.65 61.02 56.20 65.81 73.46 56.55 61.97 68.96 61.22 60.81 58.24 63.58 52.57 61.46
BGE-M3* (Chen et al., 2024) 60.09 60.60 62.37 57.34 70.69 78.97 58.78 64.12 75.60 64.72 64.61 65.31 69.85 54.20 64.80
Multilingual-E5-large* (Wang et al., 2024d) 61.91 61.97 62.91 59.40 71.30 78.08 55.21 63.41 76.53 66.55 63.75 63.67 67.32 51.55 64.54
UDEVER-Bloom-7B* (Zhang et al., 2023) 55.83 56.39 59.73 54.38 64.32 68.70 48.97 55.02 67.60 58.54 55.96 55.13 61.00 47.41 57.78
SimCSE (Gao et al., 2021a) 51.92 51.81 24.90 46.95 31.18 37.12 39.27 29.46 41.64 26.23 25.17 21.54 26.71 38.36 35.16
Contriever (Izacard et al., 2022) 49.29 44.26 26.55 44.05 33.03 39.66 38.33 32.36 45.76 26.47 23.27 22.61 22.64 39.26 34.82
GTE-large (Li et al., 2023) 62.29 51.66 33.49 50.13 38.88 44.67 43.07 30.27 51.98 27.02 20.38 22.97 22.75 41.40 38.64
BGE-en-1.5 (Xiao et al., 2023) 63.27 51.65 32.79 50.84 38.50 49.73 43.28 30.81 51.16 31.11 25.28 26.34 23.02 41.96 39.98
E5-large (Wang et al., 2024a) 60.12 52.41 26.81 51.00 37.99 39.47 43.86 31.32 53.59 28.84 24.57 23.48 22.03 43.25 38.48
ST5-XXL (Ni et al., 2021c) 58.81 60.35 44.42 58.50 41.81 24.66 53.43 25.30 52.46 15.43 18.07 17.10 21.63 38.81 37.91
GTR-XXL (Ni et al., 2021b) 58.12 54.39 41.94 53.21 37.96 24.67 50.08 25.14 53.88 15.23 17.35 15.92 22.12 40.57 36.47
E5-Mistral (Wang et al., 2024b) 66.64 61.84 61.30 59.65 58.58 72.55 58.25 54.43 66.97 62.82 56.23 55.10 47.15 50.61 59.44
@@
Table 1. Comparative analysis of model performance across multiple languages and tasks. The table presents average metrics for each model, with the highest score for each language emphasized in bold. * denotes the models trained on extensive multilingual data.
@@
SimCSE (Gao et al., 2021a) 7.41 18.35 39.71 37.95 0.18 20.72
Contriever (Izacard et al., 2022) 9.75 22.94 34.55 41.72 0.03 21.80
GTE-large (Li et al., 2023) 16.99 31.82 37.57 53.79 1.59 28.35
BGE-en-1.5 (Xiao et al., 2023) 16.64 31.19 40.40 50.77 1.11 28.02
E5-large (Wang et al., 2024a) 17.04 31.12 37.90 54.31 1.83 28.44
ST5-XXL (Ni et al., 2021c) 20.82 41.68 56.19 59.02 1.76 35.89
GTR-XXL (Ni et al., 2021b) 20.19 38.02 50.83 60.11 2.74 34.38
E5-Mistral (Wang et al., 2024b) 31.54 54.75 81.12 71.37 21.92 52.14
@@
Table 2. Cross-lingual evaluation results. The table presents average metrics for each model over all languages of the datasets, with the highest score for each language emphasized in bold.
4.1. Benchmark
Figure 2 illustrates the tasks and datasets in our benchmark. Following (Muennighoff et al., 2023), our benchmark includes five fundamental embedding tasks, with the evaluation protocol for each task adapted from the respective original papers. The benchmark involves 123 diverse datasets, including 48 Classification datasets, 24 Clustering datasets, 24 Retrieval datasets, 22 Semantic Textual Similarity STS datasets, and 5 Reranking datasets. The main metrics for each task are as follows: Classification: Accuracy, Clustering: V-measure (Rosenberg and Hirschberg, 2007), Retrieval: nDCG@10, STS: Pearson correlation based on cosine similarity (Reimers et al., 2016), and Reranking: MAP. Following (Lai et al., 2023), our benchmark covers 14 languages including 5 high-resource languages: English (en), Spanish (es), Russian (ru), French (fr), Vietnamese (vi); 6 medium-resource languages: Persian (fa), Indonesian (id), Arabic (ar), Finnish (fi), Korean (ko), Hindi (hi); 3 low-resource languages: Bengali (bn), Telugu (te), Swahili (sw).
Additionally, we evaluate models on cross-lingual retrieval tasks where the models need to perform text embedding tasks with queries and documents in different languages. These tasks feature 5 datasets, including Belebele (Bandarkar et al., 2024), MLQA (Lewis et al., 2020a), STS17, STS22 (Agirre et al., 2016), and IndicCrosslingualSTS (Ramesh et al., 2022), covering over 100 languages, including critically low-resource languages.
@@ 4.1. Benchmark
Figure 3. Performance comparison of LUSIFER and baseline models on Classification and Clustering tasks.
4.2. Experimental Setup
@@
Implementation Details. LUSIFER encompasses three key components: a multilingual encoder, a connector, and a target LLM. We employ XLM-R-large (Conneau et al., 2020) as the multilingual encoder, Mistral-7B (Jiang et al., 2023) as the English-centric target LLM, and a 2-layer feed-forward network with one trainable token as the connector. To facilitate efficient training, we leverage the LoRA framework (Hu et al., 2022) for training of LUSIFER’s components. Furthermore, we employ GradCache (Gao et al., 2021b), gradient checkpointing, mixed precision training, and FSDP (Zhao et al., 2023) to minimize GPU memory requirements. The LUSIFER architecture and its training code are built on top of the Hugging Face Transformers (Wolf et al., 2020) and Pytorch Lightning libraries (Falcon and team, 2024). We detail the training hyper-parameters for each stage in Table 3
@@
Table 3. Training hyperparameters for each stage.
Training Data. We only train LUSIFER on a diverse public English datasets. For alignment training, we use the combination of the English Wikipedia and questions-answering datasets. Specifically, we use subset of Wikitext-103 (Merity et al., 2017) and MSMARCO (Bajaj et al., 2018) for the masked reconstruction and autoregressive completion tasks, respectively. For representation finetuning, we adopt the retrieval datasets as follows: MS MARCO (Bajaj et al., 2018), NQ (Kwiatkowski et al., 2019), PAQ (Lewis et al., 2021), HotpotQA (Yang et al., 2018), SNLI (Bowman et al., 2015), SQuAD (Rajpurkar et al., 2016), ArguAna (Wachsmuth et al., 2018), FiQA (Maia et al., 2018) and FEVER (Thorne et al., 2018). To address the lack of hard negatives in these datasets, we leverage an encoder-based model (Wang et al., 2024a) to select the hard negatives on those datasets. Refer to Table 4 for the number of samples used in each dataset.
@@
Alignment Training Wikitext-103 (Merity et al., 2017) 100,000
MSMARCO (Bajaj et al., 2018) 100,000
Representation Finetuning MS MARCO (Bajaj et al., 2018) 100,000
FEVER (Thorne et al., 2018) 100,000
PAQ (Lewis et al., 2021) 100,000
SNLI (Bowman et al., 2015) 100,000
HotpotQA (Yang et al., 2018) 97,800
SQuAD (Rajpurkar et al., 2016) 97,400
FiQA (Maia et al., 2018) 6,420
NQ (Kwiatkowski et al., 2019) 3,420
ArguAna (Wachsmuth et al., 2018) 1,280